OpenStudy (anonymous):

Which of the following is the slope between the points (-2, 4) and (2, -2)? - i keep getting 2 over 4.

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(-2-4)/(2-(-2)) m=(-6)/(2+2) m=(-6)/(4) m=-3/2